OpenStudy (anonymous):

Consider the circle (x-2)^2+(y+1)^2 = 16 --> would a solution to this equation be any point where the circle crosses the x axis? Is there algebra I should be using?

OpenStudy (anonymous):

this is the equation of a circle with center (2,-1) and radius 4

OpenStudy (anonymous):

yup that's as far as I got :P

OpenStudy (anonymous):

because the statement \[(x-2)^2+(y+1)^2=16\] means the distance between (x,y) and (2,-1) is 4

OpenStudy (anonymous):

oh i see it crosses the x-axis where y = 0

OpenStudy (anonymous):

a solution to this problem is anywhere that is on the circle (x-2)^2+(y+1)^2=16 basically you have a circle with radius 4 and center (2,-1)

OpenStudy (anonymous):

so replace y by zero, solve for x

OpenStudy (anonymous):

you get \[(x-2)^2+1=16\] \[(x-2)^2=15\] \[x-2=\pm\sqrt{15}\] \[x=2\pm\sqrt{15}\]

OpenStudy (anonymous):

so the solution would be 2+1 root 15?

OpenStudy (anonymous):

2+- I MEAN

OpenStudy (anonymous):

so it crosses at two places, \[(2+\sqrt{15},0)\] and \[(2-\sqrt{15},0)\]

OpenStudy (anonymous):

yes you have it!

OpenStudy (anonymous):

cheers :)

OpenStudy (anonymous):

You rock satellite!